OpenStudy (anonymous):

Prove, in complete sentences, whether the left expression is equal to the right expression and discuss which property applies. 2(x − 3) 2x − 6 i already know that 2(x-3) =2x-6 =6/2 = x=3

OpenStudy (anonymous):

nevermind i figured it out my self

OpenStudy (radar):

Did you figure it out for sure. How did you get x=3, when there was no equation.....juest an expression. (No equal sign).!

OpenStudy (radar):

You introduced the equal sign. Originally it was a "distribution" problem to show that the expression is an identy, There are infinite number of solutions

OpenStudy (radar):

x could be 4, 2, -1 etc.

OpenStudy (anonymous):

idk but i figured it out i just needed to figure out why the expression was the same , but i still dont know what property is used there

OpenStudy (radar):

The distributive property. You distribute the 2 in the first expression getting 2x-6 which is identical to the right expression. And as you can see there are an infinite number of solutions including your result of 3.

OpenStudy (anonymous):

i understand that i just said that x = 3 because when i did 2(x-3) it turned in to 2x-6 and then i did 6/2 which is 3 but i ddnt have to do that

OpenStudy (radar):

When you substitute 3 for x in 2(x-3) you get 0.

OpenStudy (anonymous):

i know that , i substituted it when i got 2x-6

OpenStudy (radar):

Still 0 2(3)-6=0 0=0 If you had said x=5 then 2(5-3)=2(5)-6 10-6 = 10-6 4=4 You see it is an identity not a solvable equation.

OpenStudy (anonymous):

i know i realized that after i did that because it didnt make sence , and thanks for the help

OpenStudy (radar):

You are welcome, you will run into these type of problems and they can be confusing because there are no wrong answers lol.

OpenStudy (anonymous):

yeah i know

OpenStudy (anonymous):

i hate math

OpenStudy (radar):

Well that makes it even more difficult, but try to do as much "practice" exercises that you can stand.

OpenStudy (anonymous):

ok i will do , and its not that im bad at it its just too confusing